OpenStudy (anonymous):

∬▒〖(ycosx-3y2senx)dxdy i ((3π )/2〗,1) i (0,2)

ganeshie8 (ganeshie8):

\(\large \int \limits_{\frac{3\pi}{2}}^2 \int \limits_0^2 y\cos x - 3y^2 \sin x ~dx dy\)

ganeshie8 (ganeshie8):

like that ?